solve using the linear equation using elimination
3x-2y=3
-x+y=1

Answer by checkley77(12569)

3x-2y=3
-x+y=1 multiply this equation by 2 & add.
-2x+2y=2
3x-2y=3
-----------
x=5 ans.
3*5-2y=3
15-2y=3
-2y=3-15
-3y=-12
y=-12/-3
y=4 ans.
Proof:
-5+4=1
1=1